# Aligning Large Language Models for Controllable Recommendations

### v1

@@ Abstract
Inspired by the exceptional general intelligence of Large Language Models (LLMs), researchers have begun to explore their application in pioneering the next generation of recommender systems — systems that are conversational, explainable, and controllable. However, existing literature primarily concentrates on integrating domain-specific knowledge into LLMs to enhance accuracy, often neglecting the ability to follow instructions. To address this gap, we initially introduce a collection of supervised learning tasks, augmented with labels derived from a conventional recommender model, aimed at explicitly improving LLMs’ proficiency in adhering to recommendation-specific instructions. Subsequently, we develop a reinforcement learning-based alignment procedure to further strengthen LLMs’ aptitude in responding to users’ intentions and mitigating formatting errors. Through extensive experiments on two real-world datasets, our method markedly advances the capability of LLMs to comply with instructions within recommender systems, while sustaining a high level of accuracy performance.
@@
[…]
@@ 1 Introduction
Typical approaches in literature involve reformatting recommendation tasks — such as item reranking or click-through rate (CTR) prediction — into natural language constructs to facilitate the fine-tuning of LLMs. However, we have observed that LLMs fine-tuned through this straightforward method, albeit enhancing accuracy in offline evaluations, frequently generate outputs with domain-specific formatting errors. These errors may manifest as repeated items in the top-k recommendations or the inclusion of items previously interacted with by the user. Additionally, these LLMs exhibit a limited ability to adhere to user instructions. This compromises their effectiveness as viable agents in real-world recommender systems. A vidid example can be found in Section 3.4.
@@
In this paper, we investigate the alignment of an LLM for recommender systems. Our objective extends beyond merely improving the recommendation accuracy of an original LLM; we aim to significantly enhance controllability and reduce formatting errors. Drawing inspiration from the Reinforcement Learning from Human Feedback (RLHF) framework Ouyang et al. (2022), our methodology is structured into two phases: the supervised learning (SL) stage and the reinforcement learning (RL) stage. To inject domain-specific knowledge and foster recommendation-relevant control capabilities within the LLM, we devise a series of fine-tuning tasks, including item recommendation, item search, category control, and category proportion control. These tasks often necessitate generating a list of items that not only meet user instructions but also maintain high quality, despite the typically sparse ground-truth signals found in user behavior history. To tackle this issue, we propose augmenting supervised labels using predictions from traditional recommender models, such as SASRec Kang and McAuley (2018).
@@
After the SL stage, the LLM exhibits a significantly enhanced ability to follow recommendation-related instructions, surpassing existing approaches that solely fine-tune the LLM on item recommendation and search tasks. Nevertheless, the SL stage’s data generation process inherently provides only positive examples for each instruction. To address scenarios where the LLM generates suboptimal responses, we introduce an RL stage with carefully crafted reward signals to further refine the LLM’s capacity to follow instructions. To our knowledge, this is the first study that employs both SL and RL stages to align LLMs for recommendation purposes. We conduct comprehensive experiments on two real-world datasets, Steam and Amazon Movie, with results demonstrating that our method markedly improves the LLM’s ability to follow instructions while simultaneously reducing formatting errors. Our major contributions are summarized as follows:
@@
  • We introduce a novel supervised learning stage, which encompasses a suite of tasks designed for enhancing controllability and label-augmentation by a teacher recommender model, to align an LLM into an intelligent recommender agent.
@@
  • To mitigate formatting errors and improve the instruction-following ability, We further design an alignment stage based on reinforcement learning with a variety of rewards that are tailored for the nuances of the recommendation task.
@@
  • Experiment results validate that our model markedly surpasses existing LLM-based recommendation models, exhibits a robust capacity to follow users’ instructions while maintaining a high level of recommendation precision.
@@ 2.1 Intention Categories
Item-wise intention. In addition to the user profile, users may express specific desires, such that the recommended items should either exhibit particular characteristics (“I wish to watch an action movie”) or exclude them (“Please avoid suggesting any action movies”).
@@ 2.2 The SL Stage
We represent each data sample in natural language, adopting the format “Instruction: [Prompt_Content]. Output: [Response_Content]”, where [Prompt_Content] includes the detailed instruction reflecting the user’s profile and intention, and [Response_Content] contains the recommendations that fulfill the instruction. Different from traditional recommender systems that utilize item IDs, we employ item titles in both [Prompt_Content] and [Response_Content] to ensure comprehension by the LLM. Data samples are generated according to the following tasks:
@@ 2.2.1 Data Generation Tasks
Sequential Recommendation Instructions ($\bm{I_{0}}$)
@@
Category Control Instructions ($\bm{I_{1}}$)
This task corresponds to instructions with item-wise intention, which we implement into two distinct types: 1) Positive control ($\bm{I_{1}^{+C}}$), where the user hopes to receive more items in the recommendation list that match the specific category $C_{target}$. 2) Negative control ($\bm{I_{1}^{-C}}$), where users indicate a weariness towards a certain category $C_{target}$ and wish to minimize the inclusion of such items in their received recommendations as much as possible.
Category Proportion Control Instructions ($\bm{I_{2}}$)
We implement the list-wise intention into three distinct types: 1) $\bm{I_{2}^{CP\leq*}}$, in this case, the user hopes to have the proportion of the item of $C_{target}$ less than a certain value. 2) $\bm{I_{2}^{CP\approx*}}$, in this case, the user hopes that the proportion of items in $C_{target}$ will be approximately a certain value. 3) $\bm{I_{2}^{CP\geq*}}$, in this case, the user hopes that the proportion of items in $C_{target}$ is greater than a certain value.
Item Search Instructions($\bm{I_{3}}$)
To aid the LLM in memorizing item attributes, we introduce an item search task: the objective is to retrieve $k$ items belonging to a target category $C_{target}$. For this purpose, we randomly select $k$ items from $C_{target}$ to serve as the ground truth for the response.
ShareGPT($\bm{I_{4}}$)
To avoid catastrophic forgetting and preserve the general intelligence capabilities of the LLM, we follow Zeng et al. (2023) and integrate a specific proportion of ShareGPT training data into the SL stage.
@@ 2.2.2 Label Augmentation
Instructions $\bm{I_{0}}$, $\bm{I_{1}}$, and $\bm{I_{2}}$ require the inclusion of $k$ items within the response text. However, due to the typically sparse nature of user historical behavior, it is often impractical to construct a ground-truth response based solely on this information. To overcome this limitation, we employ the sequential recommender model SASRec Kang and McAuley (2018) as a teacher model to generate a set of top recommendations, $\bm{P}_{SASRec}$, for each data sample. We then curate the top-$k$ list by selecting items from $\bm{P}_{SASRec}$ that align with the given instructions. The top-$k$ list is assembled as follows: the first item is the ground-truth item (i.e., the $n^{th}$ item in the user’s history). The method for compiling the remaining $k-1$ items varies slightly: for $\bm{I_{0}}$, we fill the list with SASRec’s top predictions; for $\bm{I_{1}}$ and $\bm{I_{2}}$, we filter SASRec’s predictions to ensure the final recommendation list adheres to the specified instruction.
@@ 2.3 The RL Stage
Following previous works Touvron et al. (2023); Ouyang et al. (2022), we employ the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017) to further fine-tune the model after the SL stage. Different from Ouyang et al. (2022), scores are not produced by a reward model; instead, they are derived from reward rules that are specifically tailored for $\bm{I_{0}}$, $\bm{I_{1}}$, and $\bm{I_{2}}$. Fundamentally, rewards consist of two components: item-level rewards and list-level rewards.
@@ 2.3.1 Item-level Reward
The item-level reward assigns a score to each item generated by the LLM, serving as an immediate reward of reinforcement.
@@
Beyond the preference score, $Scores$, a control effect score, $Scores^{ctl}$, is required to gauge the extent to which an item corresponds with user intentions. Essentially, a generated item that adheres to the given instruction is awarded a high score, signifying positive reinforcement, while a non-conforming item incurs a low score, serving as negative feedback. The complete calculation is described in algorithm 1 in the Appendix. Finally, we get item-level reward $R_{item}$, which measures the overall merits of each item:
@@ 3.1.1 Dataset
We experiment with two datasets in recommender system domain: the Amazon Movies & TVs ***https://cseweb.ucsd.edu/~jmcauley/datasets/amazon˙v2/ dataset and the Steam dataset Kang and McAuley (2018). Both datasets include item category information, which aids in constructing the user’s control intentions. The statistics of our dataset are summarized in Table 1. We employed the leave-one-out approach to split the dataset. Therefore, the size of test set is consistent with the number of users in each dataset. To accelerate the validation process, we only include four types of instructions in the valid set: $I_{0}$, $I_{1}^{+C}$, $I_{1}^{-C}$, and $I_{2}^{CP\approx 50\%}$. Each type of instruction has 320 instances. Upon observing multiple metrics on the validation set, we find that SL typically converged around 30 epochs. Full details of training set on each instruction task can be found at Table 7 in the Appendix.
@@ 3.1.2 Implementation Details
We choose Llama-2-7b-chat as the foundational model for our research Touvron et al. (2023). We set model’s maximum sequence length to 1024 tokens. User behavior sequences are truncated to incorporate no more than 10 items, and item titles that are excessively lengthy are condensed to a maximum of 64 tokens. Furthermore, to accommodate the complete recommendation list within the output, we reserve a larger token count, setting the maximum output length of our model to 512 tokens. Instructions and responses are formatted using the official prompt template in Touvron et al. (2023).
We use LoRA Hu et al. (2022) to fine-tune all linear layers in Llama2-7b-chat, with trainable parameters accounting for about 0.6% of the total parameters in SL and 0.3% in RL. The optimizer is Adam. In the SL stage, the learning rate is set to 0.001, LoRA dimmention is set to 16, LoRA alpha to 8, and batch size is 64. The ShareGPT data accounted for 50% of the total training data. In the RL stage, the learning rate is $5\times 10^{-6}$. We set separate LoRA layers for actor and critic networks, with LoRA dimension at $4$ and LoRA alpha at $2$. To encourage model exploration, we set the temperature to 0.7 and the weight of entropy loss to 0.01. To ensure that the information in $R_{list}$ is not overshadowed by $R_{item}$, we amplify $R_{list}$ by a factor of ten. Source code will be released at https://github.com/microsoft/recai.
@@ 3.1.4 Baselines
We divide the compared methods into 3 classes: general LLM, fine-tuned LLM, and our method and its variants. For sequential recommendation, we additionally compared with SASRec Kang and McAuley (2018) that we used as teacher model.
@@
We use a closed-source LLM GPT-3.5-turbo-0613 https://platform.openai.com/docs/models/gpt-3-5-turbo (GPT-3.5 for short) and an open-source LLM Llama2-7b-chat Touvron et al. (2023). Due to the high cost, all GPT-3.5 results are obtained from the first $1{,}000$ samples of the test set.
@@
InstructRec Zhang et al. (2023) uses $39$ types of recommendation-related instructions constructed from user history and review data. It uses Beam search to obtain the top-k recommendation list during the test phase. Its backbone is 3B Flan-T5-XL. PALR Chen (2023) also uses Llama2-7b-chat as the backbone. We fine-tune both models using our datasets, employing data sample generation techniques as described in their respective papers. For instance, a key distinction between PALR and our supervised instruction tuning task, $I_{\{0\}}$, is that PALR does not utilize a teacher recommender model for label augmentation; instead, it relies solely on user behavior history to derive label responses.
@@
$\mathbf{Ours_{v1}}$ is a model that only uses sequential recommendation instructions ($I_{\{0\}}$) in SL. $\mathbf{Ours_{v2}}$ is a model uses all instructions ($I_{\{0,1,2,3,4\}}$) in SL. $\mathbf{Ours_{v3}}$ is a model obtained by further fine-tuning $\mathrm{Ours_{v2}}$ using RL, but without the item-level reward. $\mathbf{Ours_{full}}$ is the complete version of our method. We compare different variants of our method as an ablation study to verify the effect of different components.

### v2

@@ Abstract
Inspired by the exceptional general intelligence of Large Language Models (LLMs), researchers have begun to explore their application in pioneering the next generation of recommender systems — systems that are conversational, explainable, and controllable. However, existing literature primarily concentrates on integrating domain-specific knowledge into LLMs to enhance accuracy using a fixed task template, often overlooking the diversity of recommendation tasks and the ability of LLMs to follow recommendation-specific instructions. To address this gap, we first introduce a collection of supervised learning tasks, augmented with labels derived from a conventional recommender model, aimed at explicitly improving LLMs’ proficiency in adhering to recommendation-specific instructions. Next, we propose a reinforcement learning-based alignment procedure to enhance LLMs’ generalization ability. Extensive experiments on two real-world datasets demonstrate that our approach significantly improves the capability of LLMs to respond to instructions within recommender systems, reducing formatting errors while maintaining a high level of accuracy.
@@
Wensheng Lu1  Jianxun Lian2  Wei Zhang1  Guanghua Li1 Mingyang Zhou1  Hao Liao1∗  Xing Xie2 College of Computer Science and Software Engineering, Shenzhen University, China1 Microsoft Research Asia2 2210273060@email.szu.edu.cnjianxun.lian@outlook.com {2210275010, 2210275050}@email.szu.edu.cn {zmy, haoliao}@szu.edu.cn xingx@microsoft.com
@@ 1 Introduction
Typical approaches in the literature involve reformatting recommendation tasks — such as item reranking or click-through rate (CTR) prediction — into natural language constructs to facilitate the fine-tuning of LLMs. However, we have observed that LLMs fine-tuned through this straightforward method, albeit enhancing accuracy in offline evaluations, frequently generate outputs with domain-specific formatting errors. These errors may manifest as repeated items in the top-k recommendations or the inclusion of items previously interacted with by the user. Additionally, these LLMs exhibit a limited ability to adhere to diverse recommendation-specific instructions. This compromises their effectiveness as interactive agents in real-world recommender systems. A vivid example can be found in Section 3.4.
@@
In this paper, we investigate the alignment of an LLM for recommender systems. Our objective extends beyond merely improving the recommendation accuracy of an original LLM; we aim to significantly enhance controllability and reduce formatting errors. Drawing inspiration from the Reinforcement Learning from Human Feedback (RLHF) framework Ouyang et al. (2022), our methodology is structured into two phases: the supervised learning (SL) stage and the reinforcement learning (RL) stage. To inject domain-specific knowledge and foster recommendation-relevant control capabilities within the LLM, we devise a series of fine-tuning tasks, including item recommendation, item search, category control, and category proportion control. These tasks often necessitate generating a list of items that not only meet users’ instructions but also maintain high quality, despite the typically sparse ground-truth signals found in user behavior history. To tackle this issue, we propose augmenting supervised labels with predictions from a traditional recommender model, such as SASRec Kang and McAuley (2018). These augmented labels can help distill knowledge from the traditionally trained recommender model and meet the dynamic requirements of recommendation instructions.
@@
After the SL stage, the LLM exhibits a significantly enhanced ability to follow recommendation-related instructions, surpassing existing approaches that solely fine-tune the LLM on item recommendation and search tasks. Nevertheless, the SL stage’s data generation process inherently provides only positive examples for each instruction. To address scenarios where the LLM generates suboptimal responses, we introduce an RL stage with carefully crafted reward signals to further refine the LLM’s capacity to follow instructions. To the best of our knowledge, this is the first study that employs both SL and RL stages to align LLMs for controllable recommendation purposes. We conduct comprehensive experiments on two real-world datasets, Steam and Amazon Movie, with results demonstrating that our method markedly improves the LLM’s ability to follow instructions while simultaneously reducing formatting errors. Our major contributions are summarized as follows:
@@
  • We introduce a novel supervised learning stage, which encompasses a suite of tasks designed for enhancing controllability and label-augmentation by a teacher recommender model, to align an LLM into an interactive recommender agent.
@@
  • To mitigate formatting errors and improve the instruction-following generalization, We further design an alignment stage based on reinforcement learning with a variety of rewards that are tailored for the nuances of the controllable recommendation task.
@@
  • Experiment results validate that our model markedly surpasses existing LLM-based recommendation models, and exhibits a robust capacity to follow users’ instructions while maintaining a high level of recommendation precision. Source code is available at https://github.com/microsoft/RecAI/tree/main/RecLM-gen.
@@ 2.1 Intention Categories
Item-wise intention. In addition to the profile, users may express specific desires, such that the recommended items should either exhibit particular characteristics (“I wish to watch an action movie”) or exclude them (“Please avoid suggesting any action movies”).
@@ 2.2 The SL Stage
We represent each data sample in the recommendation task with natural language text, adopting the format “Instruction: [Prompt Content]. Output: [Response Content]”, where [Prompt Content] includes detailed instructions such as the user’s profile and intention, and [Response Content] contains the expected item recommendations that fulfill the instructions. Different from traditional recommender systems that utilize item IDs, we employ only item titles to represent items in both [Prompt Content] and [Response Content] to fully leverage LLMs’ general abilities and ensure a smooth interaction between users and our LLM-based recommender. Data samples are generated according to the following tasks:
@@ 2.2.1 Data Generation Tasks
Sequential Recommendation Instructions (${I_{0}}$)
@@
Category Control Instructions (${I_{1}}$)
This task corresponds to instructions with item-wise intention, which we implement into two distinct types: 1) Positive control (${I_{1}^{+C}}$), where the user hopes to receive more items in the recommendation list that match the specific category $C_{target}$. 2) Negative control (${I_{1}^{-C}}$), where users indicate a weariness towards a certain category $C_{target}$ and wish to reduce the inclusion of such items in their received recommendations as much as possible.
Category Proportion Control Instructions (${I_{2}}$)
We implement the list-wise intention into three distinct types: 1) ${I_{2}^{CP\leq x}}$, in this case, the user hopes to have the proportion of the item of $C_{target}$ less than a certain percentage $x$. 2) ${I_{2}^{CP\approx x}}$, in this case, the user hopes that the proportion of items in $C_{target}$ will be approximately a certain percentage. 3) ${I_{2}^{CP\geq x}}$, in this case, the user hopes that the proportion of items in $C_{target}$ is greater than a certain percentage.
Item Search Instructions (${I_{3}}$)
To aid the LLM in memorizing in-domain item attributes (in this paper, we use item category as the key attribute for illustration), we introduce an item search task: the objective is to retrieve $k$ items belonging to a target category $C_{target}$. For this purpose, we randomly select $k$ items from $C_{target}$ to serve as the ground truth for the response.
ShareGPT (${I_{4}}$)
To avoid catastrophic forgetting and preserve the general intelligence capabilities of the LLM, we follow Zeng et al. (2023) and integrate a certain proportion of ShareGPT [^1] https://huggingface.co/datasets/anon8231489123/ShareGPT˙Vicuna˙unfiltered training data into the SL stage. The ShareGPT data includes a diversity of real-world tasks in the user-assistant conversation format, which helps LLM revisit past knowledge during the in-domain fine-tuning stage.
@@ 2.2.2 Label Augmentation
Instructions ${I_{0}}$, ${I_{1}}$, and ${I_{2}}$ require the inclusion of $k$ items within the response text. However, due to the typically sparse nature of user historical behavior, it is often impractical to construct a ground-truth response based solely on this information. To overcome this limitation, we employ the sequential recommender model SASRec Kang and McAuley (2018) as a teacher model to generate a set of top recommendations, $\bm{P}_{SASRec}$, for each data sample. We then curate the top-$k$ list by selecting items from $\bm{P}_{SASRec}$ that align with the given instructions. The top-$k$ list is assembled as follows: the first item is the ground-truth item (i.e., the $n^{th}$ item in the user’s history). The method for compiling the remaining $k-1$ items varies slightly: for ${I_{0}}$, we fill the list with SASRec’s top predictions; for ${I_{1}}$ and ${I_{2}}$, we filter SASRec’s predictions to ensure the final recommendation list adhere to the specified instruction.
@@ 2.3 The RL Stage
Following previous works Touvron et al. (2023); Ouyang et al. (2022), we employ the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017) to further fine-tune the model after the SL stage. Different from Ouyang et al. (2022), scores are not produced by a reward model; instead, they are derived from reward rules that are specifically tailored for ${I_{0}}$, ${I_{1}}$, and ${I_{2}}$. Fundamentally, rewards consist of two components: item-level rewards and list-level rewards.
@@ 2.3.1 Item-level Reward
The item-level reward assigns a score to each item generated by LLM, serving as an immediate reward of reinforcement.
@@
Beyond the preference score, $Scores$, a control effect score, $Scores^{ctl}$, is required to gauge the extent to which an item corresponds with user intentions. Essentially, a generated item that adheres to the given instruction is awarded a high score, signifying positive reinforcement, while a non-conforming item incurs a low score, serving as negative feedback. The complete calculation is described in algorithm 1 in the Appendix.
Finally, we get item-level reward $R_{item}$, which measures the overall merits of each item:
@@ 2.3.2 List-level Reward
2.3.3 RL Implementation Notes
We follow the Transformer Reinforcement Learning [^2] https://github.com/huggingface/trl package to implement the reinforcement learning stage. We set up LoRA layers Hu et al. (2022) for both the policy network and the critic network, with a LoRA rank of 4 and a LoRA alpha of 2. During the RL sampling phase, we sample 2 responses with a temperature of $0.7$ for each instruction. The final reward is calculated by combining item-level reward, list-level reward, and a KL penalty:
$\bm{R}_{\mathrm{final}}[\bm{y}]=\bm{R}_{item}[\bm{y}]+\bm{R}_{list}[\bm{y}]-\eta\bm{\mathrm{KL}}(\pi_{\theta}^{\mathrm{RL}},\pi^{\mathrm{SFT}})[\bm{y}]$ (6)
where $\eta$ is set to $0.3$. $\bm{y}$ represents a generated response (which is a token sequence) for an instruction sample. $\bm{\mathrm{KL}}[\bm{y}]$ represents there is a KL penalty on each token in $\bm{y}$. $\bm{R}_{item}[\bm{y}]$ means there is an item-level reward at the last token of each item title. $\bm{R}_{list}[\bm{y}]$ represents that there is a list-level reward on the ending token of $\bm{y}$. To ensure that the information in $R_{list}$ is not overshadowed by $R_{item}$, we amplify $R_{list}$ by a factor of 10. Additionally, we employ reward whitening to enhance the stability of training. We use Generalized Advantage Estimation (GAE) to estimate the advantage values, with hyperparameters $\gamma=0.99,\lambda=0.95$. During the loss calculation phase, we set the clipping range for the probability ratio to $[1-\epsilon,1+\epsilon]$, where $\epsilon=0.2$. The loss weight for the critic network is $0.5$.
@@ 3.1.1 Dataset
We experiment with two popular datasets in the recommender system domain: the Amazon Movies & TVs  [^3] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon˙v2/ dataset (Movie for short) and the Steam dataset Kang and McAuley (2018). Both datasets include item category information, which aids in constructing the user’s control intentions. Basic statistics of datasets are summarized in Table 1. We employ the leave-one-out approach Kang and McAuley (2018) to split the dataset. Therefore, the size of test set is consistent with the number of users in each dataset. To accelerate the validation process, we only include four types of instructions in the valid set: $I_{0}$, $I_{1}^{+C}$, $I_{1}^{-C}$, and $I_{2}^{CP\approx 50\%}$, and each type of instruction has 320 instances. Upon observing multiple metrics on the validation set, we find that SL typically converged around 30 epochs. Full details of the training set on each instruction task can be found at Table 8 in the Appendix.
@@ 3.1.2 Implementation Details
We choose Llama-2-7b-chat as the foundational model for our research Touvron et al. (2023). We set the model’s maximum sequence length to 1024 tokens. User behavior sequences are truncated to incorporate no more than 10 items, and excessively lengthy item titles are condensed to a maximum of 64 tokens. Furthermore, to accommodate the complete recommendation list within the output, we reserve a larger token count, setting the maximum output length of our model to 512 tokens. Instructions and responses are formatted using the official prompt template in Touvron et al. (2023).
We use LoRA Hu et al. (2022) to fine-tune all linear layers in Llama2-7b-chat, with trainable parameters accounting for about 0.6% of the total parameters in SL and 0.3% in RL. The optimizer is Adam. In the SL stage, the learning rate is set to $0.001$, the LoRA dimension is set to 16, the LoRA alpha to 8, and the batch size is 64. The ShareGPT data accounted for 50% of the total training data. In the RL stage, the learning rate is $5\times 10^{-6}$. We set separate LoRA layers for actor and critic networks, with LoRA dimension at $4$ and LoRA alpha at $2$. To encourage model exploration, we set the temperature to $0.7$ and the weight of entropy loss to 0.01. We release the source code at: https://github.com/microsoft/RecAI/tree/main/RecLM-gen.
@@ 3.1.4 Baselines
We divide the compared methods into 3 classes: general LLM, fine-tuned LLM, and our method and its variants. For sequential recommendation, we additionally compared with SASRec Kang and McAuley (2018) that we used as the teacher model.
@@
We use a closed-source LLM GPT-3.5-turbo-0613  [^4] https://platform.openai.com/docs/models/gpt-3-5-turbo (GPT-3.5 for short) and an open-source LLM Llama2-7b-chat Touvron et al. (2023). Due to the high cost, all GPT-3.5 results are obtained from the first $1{,}000$ samples of the test set.
@@
InstructRec Zhang et al. (2023) uses $39$ types of recommendation-related instructions constructed from user history and review data. It uses Beam search to obtain the top-k recommendation list during the test phase. Its backbone is 3B Flan-T5-XL. PALR Chen (2023) also uses Llama2-7b-chat as the backbone. We fine-tune both models using our datasets, employing data sample generation techniques as described in their respective papers. For instance, a key distinction between PALR and our supervised instruction tuning task, $I_{0}$, is that PALR does not utilize a teacher recommender model for label augmentation; instead, it relies solely on user behavior history to derive label responses.
@@
$\mathbf{Ours_{v1}}$ is a model that only uses sequential recommendation instructions ($I_{0}$) in SL. $\mathbf{Ours_{v2}}$ is a model uses all instructions ($I_{\{0,1,2,3,4\}}$) in SL. $\mathbf{Ours_{v3}}$ is a model obtained by further fine-tuning $\mathrm{Ours_{v2}}$ with RL, but without the item-level reward. $\mathbf{Ours_{full}}$ is the complete version of our method. We compare different variants of our method as an ablation study to verify the effect of different components.